OpenStudy (anonymous):

Find the area of the region bounded by the functions f(x) = x4 and g(x) =sq root IxI . (2 points) 7/15 14/15 28/15 None of these

OpenStudy (solomonzelman):

I won't do your problem, but I will do a similar one. \(\color{#b0000b}{ \displaystyle \small \bf\text{----------------------------------------------------------------} }\) Find the area of the region bounded by \(\color{#000000}{ f(x)=\displaystyle x^2 }\) and \(\color{#000000}{ \displaystyle g(x)=\sqrt{x~} }\). \(\color{#b0000b}{ \displaystyle \small \bf\text{----------------------------------------------------------------} }\) At first, the sketch of the region (shaded in black). ` https://www.desmos.com/calculator/rrfdhupsgh ` (copy paste this link to a new tab, and look at it) You will see that your region starts from \(\color{#000000}{ x=0 }\) and ends at \(\color{#000000}{ x=1 }\), and from there you know the limits of integration. Note, that if you were to just write \(\color{#000000}{ \displaystyle\int\limits_0^1\sqrt{{\tiny~}x{\tiny~}}{\tiny~}dx }\) then you are finding the following area: ` https://www.desmos.com/calculator/r2yztbqshj ` and that clearly exceeds the bounded region! What if you tried to write \(\color{#000000}{ \displaystyle\int\limits_0^1x^2{\tiny~}dx }\) then you are finding the following area: ` https://www.desmos.com/calculator/zl97xehgig ` and that is just the area between the region and the x-axis. If you look at the graphs tho, you will not that you can find the area of your region \(\Omega\), by subtracting these two areas: \(\color{#000000}{ \displaystyle\Omega=\left(\int\limits_0^1\sqrt{{\tiny~}x{\tiny~}}{\tiny~}dx\right)-\left(\int\limits_0^1x^2{\tiny~}dx\right) }\) Look carefully at the graph below and make sure you understand this observation: ` https://www.desmos.com/calculator/jubjyhaijj ` Now, after conceptualizing everything, we can proceed to doing the math. The integrals (like limits and derivatives) are linear (if they have the same limits of integration), and that implies, \(\color{#000000}{ \displaystyle\Omega=\int\limits_0^1\left(\sqrt{{\tiny~}x{\tiny~}}-x^2\right) {\tiny~}dx }\) So, this is the integral you need to setup. Now, we can apply the power rule term-by-term and plug in the limits of integration. \(\color{#000000}{ \displaystyle\Omega=\int\limits_0^1\left.\left(x^{1/2}-x^2\right) {\tiny~}dx=\left(\frac{2}{3}x^{3/2}-\frac{1}{3}x^3\right)\right|_{x=0}^{x=1} }\) At \(x=0\) this is obviously 0, so we can write, \(\color{#000000}{ \displaystyle\Omega=\left.\left(\frac{2}{3}x^{3/2}-\frac{1}{3}x^3\right)\right|^{x=1} =\frac{2}{3}-\frac{1}{3}=\color{red}{\frac{1}{3}} }\)

OpenStudy (solomonzelman):